OpenStudy (anonymous):

Simplify each expression. Rationalize all denominators. Assume that all variables are positive. (2-√5)(2+√5)

OpenStudy (2bornot2b):

4-5=-1

OpenStudy (hoblos):

(a-b)(a+b) = a^2 - b^2 (2-√5)(2+√5) = 4-5 =-1